OpenStudy (anonymous):

Write an equation for the nth term of the given sequence. 8,72,648,...,a

hartnn (hartnn):

notice : 72 =8*9 648 = 72*9 so, next term is 9 times the previous. ---> common ratio =r= 9 1st term = a1 = 8 Then the n'th term of given geometric progression will be \[\large a_n=a_1.r^{n-1}\] just put r and a1 values and you get your equation :)